OpenStudy (anonymous):

The local radio station is giving away some tickets to the circus.Adult tickets cost the station $ 8 each while children's tickets cost $6. The station wants to spend no more than $172. Let x = the number of adult tickets and y= the number of children's tickets. show all possible solutions by graphing a linear inequality.

Parth (parthkohli):

I have some hints in store. - The money that they spend will be the \(\textbf{sum}\) of the money they spend on children and adults. - The money that they spend on guys will be \(\text{number of guys} \times \text{money for 1 guy}\). - The inequality sign that you use is less than or equal to/\(\le\)

OpenStudy (anonymous):

thank. how to graphing

Parth (parthkohli):

http://wolframalpha.com will help. Note that I am not the best teacher when it comes to graphing... :(

OpenStudy (sasogeek):

x=number of adult tickets adult tickets cost $8 hence total amount in $ of adult tickets = $8x y=number of children tickets children tickets cost $6 hence total amount in $ of children tickets = $6y the total of these tickets if all are sold should be equal to or less that how much is on budget that the radio station wants to spend, aka $172. hence, 8x + 6y \(\le \) 172

OpenStudy (sasogeek):

does the above logic make any sense?

OpenStudy (anonymous):

Thank you for answer

OpenStudy (sasogeek):

can you graph that?